# Spectral properties of symmetrized AMV operators

### v2

@@ 3.1. Definitions
For this subsection, we consider a metric measure space $(X,\mathsf{d},\mu)$ satisfying ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$) for some fixed $r>0$.
@@
$\|\tilde{\Delta}_{r}\|_{2\to 2}\leq\frac{1}{r^{2}}\left(\|V(\cdot,r)^{-1}\|_{1}^{1/2}+c(r)^{1/2}+c(r)+1\right).$ (25)
@@
The self-adjointness of $\tilde{\Delta}_{r}$ is obvious because $\tilde{A}_{r}$ and $[\tilde{A}_{r}1]\mathrm{I}$ are self-adjoint too. The boundedness is a consequence of Lemma 2.10. Indeed,
@@
$\leq\frac{1}{r^{2}}\left(\frac{\|V(\cdot,r)^{-1}\|_{1}^{1/2}+c(r)^{1/2}}{2}+\|\tilde{A}_{r}1\|_{\infty}\underbrace{\|I\|_{2\to 2}}_{=1}\right)$
$\leq\frac{1}{r^{2}}\left(\frac{\|V(\cdot,r)^{-1}\|_{1}^{1/2}+c(r)^{1/2}}{2}+\frac{\|A_{r}1\|_{\infty}+\|A_{r}^{*}1\|_{\infty}}{2}\right)$
$=\frac{1}{r^{2}}\left(\frac{\|V(\cdot,r)^{-1}\|_{1}^{1/2}+c(r)^{1/2}}{2}+\frac{1+c(r)}{2}\right)$
@@
×$\qquad\qquad\qquad\qquad\times\frac{(f(x)-f(y))(\psi(x)-\psi(y))}{r^{2}}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x).$
@@
Observe that (26) implies that $-\tilde{\Delta}_{r}$ is a non-negative operator, since for any $f\in L^{2}(X,\mu)$,
@@
We let $\sigma(-\tilde{\Delta}_{r})$ denote the spectrum of $-\tilde{\Delta}_{r}$, that is to say, the set of elements $\lambda\in\mathbb{C}^{*}$ such that $-\tilde{\Delta}_{r}-\lambda\mathrm{I}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is not a bijection.
@@
  • $\sigma_{\color[rgb]{.75,0,.25}\text{c}}(T)$ is the compression spectrum, that is to say, the set of $\lambda\in\mathbb{C}^{+}$ whose conjugate $\bar{\lambda}$ is an eigenvalue of the adjoint $T^{*}$,
@@
Since $-\tilde{\Delta}_{r}$ is self-adjoint and non-negative, we know that
@@
Let $\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$ denote the essential spectrum of $-\tilde{\Delta}_{r}$, i.e. the closed subset of $\sigma(-\tilde{\Delta}_{r})$ made of those $\lambda$ such that $-\tilde{\Delta}_{r}-\lambda\mathrm{I}$ is not a Fredholm operator. Since $-\tilde{\Delta}_{r}$ is self-adjoint, the Fischer–Polyà minimum-maximum principle (see e.g. [WS72, p.12]) asserts that if there exists a positive integer $N$ such that $\tilde{\lambda}_{N,r}<\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$ then $-\tilde{\Delta}_{r}$ admits $N+1$ isolated eigenvalues
@@ 3.2. Spectral properties
Our first spectral result on $\tilde{\Delta}_{r}$ is the following.
@@
Let $(X,\mathsf{d},\mu)$ be a metric measure space satisfying ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$) for some fixed $r>0$. Assume that $\lambda\in\sigma(-\tilde{\Delta}_{r})$ satisfies
@@
Then $\lambda$ is an isolated eigenvalue of $-\tilde{\Delta}_{r}$ which does not belong to $\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$.
@@
Thus $f_{n}$ converges in $L^{2}(X,\mu)$ to the limit function $f$. Using continuity of $-\tilde{\Delta}_{r}$ we conclude that
$-\tilde{\Delta}_{r}f=\lambda f,$ (33)
@@
$-\tilde{\Delta}_{r}f_{n}=\lambda_{n}f_{n}.$ (35)
@@
$-\frac{\tilde{A}_{r}}{r^{2}}f_{n}=(\lambda_{n}-\frac{[\tilde{A}_{r}1]}{r^{2}})f_{n}.$ (36)
@@
Let us now prove that $-\tilde{\Delta}_{r}-\lambda I$ is a Fredholm operator.
@@
$-\frac{\tilde{A}_{r}f_{n}}{r^{2}}=\left(\lambda-\frac{[\tilde{A}_{r}1]}{r^{2}}\right)f_{n}.$ (40)
@@
To show that the image of $-\tilde{\Delta}_{r}-\lambda I$ is closed, consider a sequence $g_{n}:=(-\tilde{\Delta}_{r}-\lambda I)(f_{n})$ such that $g_{n}\rightarrow g$. Similarly to before, we can conclude that since $g_{n}$ converges, then $f_{n}$ converges to some $f$, and thus $g=(-\tilde{\Delta}_{r}-\lambda I)(f)$. ∎
@@
Let $(X,\mathsf{d},\mu)$ be a metric measure space such that for some $r_{0}>0$ both ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$) hold for any $r\in(0,r_{0})$. Then
@@
Let $(X,\mathsf{d},\mu)$ be a connected metric measure space satisfying ($\mathrm{A}_{r}$) and ($\mathrm{A}_{r}^{*}$) for some fixed $r>0$. Then the kernel of $\tilde{\Delta}_{r}$ contains constant functions only, and $\tilde{E}_{r}$ defines a scalar product on
@@
[…]
@@
[…]
@@
$\mu(X\backslash A)=0,$ (46)
Since the measure of any open ball is positive, $X\backslash A$ cannot contain such a ball. This implies that $A$ is a dense subset of $X$. Take $z\in A$ and let $c=f(z)$. Consider
@@
Since $V$ and $W$ are open sets whose union contains $A$ which is dense in $X$, we must have
$W\cup V=X.$ (49)
Since $W$ and $V$ form an open cover of $X$, and $X$ is connected, if both $V$ and $W$ are different from the empty set, then there exist $x\in I$ and $y\in I^{\prime}$ such that
@@
This is not possible since $f(x)\neq f(y)$ and $\mu(B_{r}(x)\cap B_{r}(y))>0$. This implies that $I=A$ and that $f(w)=c$ for $\mu$-a.e. $w\in X$. Then $-\tilde{\Delta}_{r}$ has a non-trivial kernel consisting of the constant functions only. Moreover, since $-\tilde{\Delta}_{r}$ is non-negative (Remark 3.5), we get the desired property on (42).
We are now in a position to prove Theorem 1. We recall that the context of this statement is a compact uniformly locally doubling metric measure space $(X,\mathsf{d},\mu)$.
@@
Let $r_{0},C$ be the parameters of the local doubling property of $(X,\mathsf{d},\mu)$, see Definition 2.1. Then for any $i\in\{0,\ldots,k\}$ and $r\in(0,r_{0})$,
@@
By Corollary 3.8, there exists $r_{k}\in(0,\tilde{r}_{k})$ such that $\tilde{C}<\min\sigma_{\mathrm{ess}}(-\tilde{\Delta}_{r})$ for any $r\in(0,r_{k})$. Then Remark 3.6 implies that for such an $r$ the operator $-\tilde{\Delta}_{r}$ admits $k+1$ eigenvalues $\lambda_{0}(-\tilde{\Delta}_{r})\leq\lambda_{1}(-\tilde{\Delta}_{r})\leq\ldots\leq\lambda_{k}(-\tilde{\Delta}_{r})$ such that $\lambda_{i}(-\tilde{\Delta}_{r})=\tilde{\lambda}_{i,r}$ for any $i\in\{0,\ldots,k\}$. That $\lambda_{0}(-\tilde{\Delta}_{r})=0$ follows from Proposition 3.9. Moreover, by Remark 3.6 and Proposition 3.9, we know that
@@
where $\Pi(X,\mu)$ is as in (42). Since $-\Delta_{r}$ has a kernel which is $L^{2}$-orthogonal to $\Pi(X,\mu)$ (Proposition 3.9), we have $\tilde{E}_{r}(f)>0$ for any $f\in\Pi(X,\mu)$, hence we get
@@ 4.1. A preliminary lemma
$\operatorname{sinc}(\rho)\coloneqq\begin{cases}\displaystyle\frac{\sin(\pi\rho)}{\pi\rho}\qquad&\rho\in\mathbb{R}\backslash\{0\},\\ 1&\rho=0,\end{cases}$
@@
If (56) were to fail, due to the previous line, there would exist sequences $(r_{n})\subset(0,+\infty)$ and $(p^{(n)})\subset\mathbb{Z}^{m}\backslash\{0\}$ such that $r_{n}\to 0$ and
@@ 4.2. Torus
$\tilde{Q}_{r}(x)=\pi(Q_{r}(\tilde{x}))$ (67)
where $\tilde{x}$ is any element in $\pi^{-1}(x)$, and $Q_{r}(\tilde{x})$ is as in (65). Then
@@
$-\tilde{\Delta}_{r_{n}}f_{n}=\lambda_{1}(-\tilde{\Delta}_{r_{n}})f_{n},\quad\quad\lambda_{1}(-\tilde{\Delta}_{r_{n}})\rightarrow 0.$ (70)
@@
We consider the Fourier decomposition of $f_{n}$, $1_{\tilde{Q}_{r_{n}}(0)}$, and $-\tilde{\Delta}_{r_{n}}f_{n}$, namely
@@
$-\tilde{\Delta}_{r_{n}}f_{n}=\sum_{p\in\mathbb{Z}^{m}}c_{p,n}e_{p}$
@@
Using Lemma 4.1, for $\0p\in\mathbb{Z}\backslash\{0\}$ we conclude that there exists $\alpha>0$ such that
@@
$\|-\tilde{\Delta}_{r_{n}}f_{n}\|_{L^{2}(\mathbb{T}^{m})}^{2}=\sum_{p\in\mathbb{Z}^{m}}|c_{p,n}|^{2}\geq\sum_{p\in\mathbb{Z}^{m}\backslash\{0\}}|a_{p,n}|^{2}\alpha^{2}=\alpha^{2}\|f_{n}\|_{L^{2}(\mathbb{T}^{m})}^{2},$ (77)
@@ 4.3. Shrinking hypercubes
Let $C_{1},\ldots,C_{2^{m}}$ denote the $2^{m}$ sets of the form $I_{1}\times\ldots\times I_{m}$ where each $I_{i}$ is either $[-1,0]$ or $[0,1]$. For any $j\in\{1,\ldots,2^{m}\}$, set
@@
where $\varepsilon_{i}$ is $1$ if $I_{i}=[0,1]$ and $-1$ otherwise. Note that $N_{j}$ is an isometry which preserves the Lebesgue measure, and that $T\circ\pi\circ N_{j}$ is equal to the identity. Then
@@
$0<\lambda_{1}(-\tilde{\Delta}_{\bar{r}_{n},\mathfrak{T}^{m}})\leq 2^{m}\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)})\to 0$
@@
×$\qquad\qquad\quad\times\left(\frac{f_{n}(x)-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$
@@
$\pi_{i}(R_{n,k}(x))=\begin{cases}[0,\overline{r}_{n}+|\overline{x}_{i}|]&\text{ if }(1)\\ [-\overline{r}_{n}+|\overline{x}_{i}|,1]&\text{ if }(2)\\ [-\overline{r}_{n}+|\overline{x}_{i}|,1]&\text{ if }(3)\\ [-\overline{r}_{n},\overline{r}_{n}]&\text{ if }(4)\end{cases}$ (87)
$\pi_{l}(R_{n,k}(x))=\begin{cases}\pi_{i}(R_{n,l}(x))=\pi_{i}(R_{n,k}(x))&\text{ if }\pi_{i}(C_{j})=\pi_{i}(C_{k})\\ [0,\overline{r}_{n}-|\overline{x}_{i}|]&\text{ if }(1)\text{ and }\neg\left(\pi_{i}(C_{j})=\pi_{i}(C_{k})\right)\\ [2-\overline{r}_{n}-|\overline{x}_{i}|,1]&\text{ if }(2)\text{ and }\neg\left(\pi_{i}(C_{j})=\pi_{i}(C_{k})\right)\\ [2-\overline{r}_{n}-|\overline{x}_{i}|,1]&\text{ if }(3)\text{ and }\neg\left(\pi_{i}(C_{j})=\pi_{i}(C_{k})\right)\\ \emptyset&\text{ if }(4)\text{ and }\neg\left(\pi_{i}(C_{j})=\pi_{i}(C_{k})\right).\end{cases}$ (88)
@@
×$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(y))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)$ (91)
@@
×$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(y))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)$ (93)
@@
×$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(\pi(y)))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (95)
@@
×$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(\pi(N_{j}(y))))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y).$ (97)
@@
×$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (100)
@@
×$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (102)
@@
×$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (104)

### v3

@@ 3.1. Definitions
For this subsection, we consider a metric measure space $(X,\mathsf{d},\mu)$ satisfying ($\mathrm{I}_{r}$) for some fixed $r>0$.
@@
$\|\tilde{\Delta}_{r}\|_{2\to 2}\leq\frac{1}{2r^{2}}\left(2\|A_{r}^{*}1\|_{\infty}^{1/2}+\|A_{r}^{*}1\|_{\infty}+1\right).$ (25)
@@
The self-adjointness of $\tilde{\Delta}_{r}$ is obvious because $\tilde{A}_{r}$ and $\tilde{A}_{r}1]\mathrm{I}$ are self-adjoint too. The boundedness is a consequence of Lemma 2.12. Indeed,
@@
$\leq\frac{1}{r^{2}}\left(\|A_{r}^{*}1\|_{\infty}^{1/2}+\|\tilde{A}_{r}1\|_{\infty}\underbrace{\|I\|_{2\to 2}}_{=1}\right)$
$\leq\frac{1}{r^{2}}\left(\|A_{r}^{*}1\|_{\infty}^{1/2}+\frac{\|A_{r}1\|_{\infty}+\|A_{r}^{*}1\|_{\infty}}{2}\right)$
$=\frac{1}{r^{2}}\left(\|A_{r}^{*}1\|_{\infty}^{1/2}+\frac{1+\|A_{r}^{*}1\|_{\infty}}{2}\right)$
@@
$\qquad\qquad\qquad\qquad\times\frac{(f(x)-f(y))(\psi(x)-\psi(y))}{r^{2}}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x).$
Remark 3.5.
A suitable use of the Fubini–Tonelli theorem shows that the energy functional $\tilde{E}_{r}(f)$ equals the approximate Korevaar–Schoen energy [KS93]
$\frac{1}{2}\int_{X}\fint_{B_{r}(x)}\frac{|f(y)-f(x)|^{2}}{r^{2}}\mathop{}\!\mathrm{d}\mu(y)\mathop{}\!\mathrm{d}\mu(x).$
@@
Observe that (26) implies that $\tilde{\Delta}_{r}$ is a non-negative operator, since for any $f\in L^{2}(X,\mu)$,
@@
We let $\sigma(-\tilde{\Delta}_{r})$ denote the spectrum of $\tilde{\Delta}_{r}$, that is to say, the set of elements $\lambda\in\mathbb{C}^{*}$ such that $\tilde{\Delta}_{r}-\lambda\mathrm{I}:L^{2}(X,\mu)\to L^{2}(X,\mu)$ is not a bijection.
@@
  • $\sigma_{\text{c}}(T)$ is the compression spectrum, that is to say, the set of $\lambda\in\mathbb{C}^{+}$ whose conjugate $\bar{\lambda}$ is an eigenvalue of the adjoint $T^{*}$,
@@
Since $\tilde{\Delta}_{r}$ is self-adjoint and non-negative, we know that
@@
Let $\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$ denote the essential spectrum of $\tilde{\Delta}_{r}$, i.e. the closed subset of $\sigma(-\tilde{\Delta}_{r})$ made of those $\lambda$ such that $\tilde{\Delta}_{r}-\lambda\mathrm{I}$ is not a Fredholm operator. Since $\tilde{\Delta}_{r}$ is self-adjoint, the Fischer–Polyà minimum-maximum principle (see e.g. [WS72, p.12]) asserts that if there exists a positive integer $N$ such that $\tilde{\lambda}_{N,r}<\min\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$ then $\tilde{\Delta}_{r}$ admits $N+1$ isolated eigenvalues
@@ 3.2. Spectral properties
Our first spectral result on $\tilde{\Delta}_{r}$ is the following. Note that we need the compactness of $\tilde{A}_{r}$ here, thus we assume ($\mathrm{II}_{r}$).
@@
Let $(X,\mathsf{d},\mu)$ be a metric measure space satisfying ($\mathrm{II}_{r}$) for some fixed $r>0$. Assume that $\lambda\in\sigma(-\tilde{\Delta}_{r})$ satisfies
@@
Then $\lambda$ is an isolated eigenvalue of $\tilde{\Delta}_{r}$ which does not belong to $\sigma_{\text{ess}}(-\tilde{\Delta}_{r})$.
@@
Thus $f_{n}$ converges in $L^{2}(X,\mu)$ to the limit function $f$. Using continuity of $\tilde{\Delta}_{r}$ we conclude that
$\tilde{\Delta}_{r}f=\lambda f,$ (33)
@@
$\tilde{\Delta}_{r}f_{n}=\lambda_{n}f_{n}.$ (35)
@@
$\frac{\tilde{A}_{r}}{r^{2}}f_{n}=(\lambda_{n}-\frac{[\tilde{A}_{r}1]}{r^{2}})f_{n}.$ (36)
@@
Let us now prove that $\tilde{\Delta}_{r}-\lambda I$ is a Fredholm operator.
@@
$\frac{\tilde{A}_{r}f_{n}}{r^{2}}=\left(\lambda-\frac{[\tilde{A}_{r}1]}{r^{2}}\right)f_{n}.$ (40)
@@
To show that the image of $\tilde{\Delta}_{r}-\lambda I$ is closed, consider a sequence $g_{n}:=(-\tilde{\Delta}_{r}-\lambda I)(f_{n})$ such that $g_{n}\rightarrow g$. Similarly to before, we can conclude that since $g_{n}$ converges, then $f_{n}$ converges to some $f$, and thus $g=(-\tilde{\Delta}_{r}-\lambda I)(f)$. ∎
@@
Let $(X,\mathsf{d},\mu)$ be a metric measure space such that for some $r_{0}>0$ the assumption ($\mathrm{II}_{r}$) holds for any $r\in(0,r_{0})$. Then
@@
Let $(X,\mathsf{d},\mu)$ be a connected metric measure space satisfying ($\mathrm{II}_{r}$) for some fixed $r>0$. Then the kernel of $\tilde{\Delta}_{r}$ contains constant functions only, and $\tilde{E}_{r}$ defines a scalar product on
@@
$\int_{X}1_{B_{r}(x)}(y)\left(\frac{1}{V(x,r)}+\frac{1}{V(y,r)}\right)\left(\frac{f(x)-f(y)}{r}\right)^{2}d\mu(y)=0$ (43)
@@
Consider $F\coloneqq\{x\in X:\text{$f(x)$ is a well-defined real number}\}$ and
@@
[…]
Take $z\in A$ and let $c=f(z)$. Consider
@@
Since $V$ and $W$ are open sets whose union contains $A$ which has full measure in $X$, we must have
$W\cup V=X,$ (49)
[…]
@@
This is not possible since $f(x)\neq f(y)$ and $\mu(B_{r}(x)\cap B_{r}(y))>0$. This implies that $I=A$ and that $f(w)=c$ for $\mu$-a.e. $w\in X$. Then $\tilde{\Delta}_{r}$ has a non-trivial kernel consisting of the constant functions only. Moreover, since $\tilde{\Delta}_{r}$ is non-negative (Remark 3.8), we get the desired property on (42). ∎
We are now in a position to prove Theorem 1. We recall that the context of this statement is a compact uniformly locally doubling metric measure space $(X,\mathsf{d},\mu)$. The compactness of the space ensures that ($\mathrm{II}_{r}$) holds for any $r>0$, see Lemma 2.6.
@@
Let $r_{0},C$ be the parameters of the uniform local doubling property of $(X,\mathsf{d},\mu)$, see Definition 2.8. Then for any $i\in\{0,\ldots,k\}$ and $r\in(0,r_{0})$,
@@
By Corollary 3.13, there exists $r_{k}\in(0,\tilde{r}_{k})$ such that $\tilde{C}<\min\sigma_{\mathrm{ess}}(-\tilde{\Delta}_{r})$ for any $r\in(0,r_{k})$. Then Remark 3.11 implies that for such an $r$ the operator $\tilde{\Delta}_{r}$ admits $k+1$ eigenvalues $\lambda_{0}(-\tilde{\Delta}_{r})\leq\lambda_{1}(-\tilde{\Delta}_{r})\leq\ldots\leq\lambda_{k}(-\tilde{\Delta}_{r})$ such that $\lambda_{i}(-\tilde{\Delta}_{r})=\tilde{\lambda}_{i,r}$ for any $i\in\{0,\ldots,k\}$. That $\lambda_{0}(-\tilde{\Delta}_{r})=0$ follows from Proposition 3.14. Moreover, by Remark 3.11 and Proposition 3.14, we know that
@@
where $\Pi(X,\mu)$ is as in (42). Since $\Delta_{r}$ has a kernel which is $L^{2}$-orthogonal to $\Pi(X,\mu)$ (Proposition 3.14), we have $\tilde{E}_{r}(f)>0$ for any $f\in\Pi(X,\mu)$, hence we get
@@ 4.1. A preliminary lemma
[…]
@@
[…]
@@ 4.2. Torus
$)\tilde{Q}_{r}(x)=\pi(Q_{r}(\tilde{x}$)) (67)
[…]
@@
$\tilde{\Delta}_{r_{n}}f_{n}=\lambda_{1}(-\tilde{\Delta}_{r_{n}})f_{n},\quad\quad\lambda_{1}(-\tilde{\Delta}_{r_{n}})\rightarrow 0.$ (70)
@@
We consider the Fourier decomposition of $f_{n}$, $1_{\tilde{Q}_{r_{n}}(0)}$, and $\tilde{\Delta}_{r_{n}}f_{n}$, namely
@@
$\tilde{\Delta}_{r_{n}}f_{n}=\sum_{p\in\mathbb{Z}^{m}}c_{p,n}e_{p}$
@@
[…]
@@
[…]
@@ 4.3. Shrinking hypercubes
Let $C_{1},\ldots,C_{2^{m}}$ denote the $2^{m}$ sets of the form $I_{1}\times\ldots\times I_{m}$ where each $I_{i}$ is either [1,0][-1,0] or $[0,1]$. For any $j\in\{1,\ldots,2^{m}\}$, set
@@
where $\varepsilon_{i}$ is $1$ if $I_{i}=[0,1]$ and 1-1 otherwise. Note that $N_{j}$ is an isometry which preserves the Lebesgue measure, and that $T\circ\pi\circ N_{j}$ is equal to the identity. Then
@@
$0<\lambda_{1}(-\tilde{\Delta}_{\bar{r}_{n},\mathfrak{T}^{m}})\leq 2^{m}\lambda_{1}(-\tilde{\Delta}_{\overline{r}_{n},\mathfrak{Q}^{m}(1)})\to 0.$
@@
$\qquad\qquad\quad\times\left(\frac{f_{n}(x)-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$
@@
[…]
@@
$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(y))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)$ (91)
@@
$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(y))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathbb{L}^{m}(y)$ (93)
@@
$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(\pi(y)))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (95)
@@
$\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(T(\pi(N_{j}(y))))}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y).$ (97)
@@
$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (100)
@@
$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (102)
@@
$\quad\quad\quad\quad\times\left(\frac{f_{n}(T(x))-f_{n}(y)}{\overline{r}_{n}}\right)^{2}\mathop{}\!\mathrm{d}\mathcal{L}^{m}(y)$ (104)